# A priori error analysis of the proximal Galerkin method

### v1

@@ 3.4. Best approximation error estimates
$\leq\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle+\langle E^{\prime}(\overline{u}_{h}^{\ell})-E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle$
$\leq\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle+M\|u-\overline{u}_{h}^{\ell}\|_{V}\|u-\mathcal{U}_{h}(\psi)\|_{V}$
$\leq\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle+\frac{\nu}{2}\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}+\frac{M^{2}}{2\nu}\|u-\mathcal{U}_{h}(\psi)\|_{V}^{2}\,.$
Canceling terms, we arrive at the following estimate:
(74) $E(u_{h}^{\ell})-E(u)\leq\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(u),\mathcal{U}_{h}(\psi)-u\rangle+\frac{M^{2}}{2\nu}\|u-\mathcal{U}_{h}(\psi)\|_{V}^{2}\,,$
which allows us to obtain an estimate of the error in the iterates.
At this stage, we set $u=u^{\ast}$ and let $v\in K$. We then use 70 and 9 to deduce that
@@
Thus, adding $\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}\rangle$ to both sides of 74, we find that
(75) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\\ \leq\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|u^{\ast}-\mathcal{U}_{h}(\psi)\|_{V}^{2}+\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}+\mathcal{U}_{h}(\psi)-u^{\ast}\rangle\,.$
The result follows because the choices of $\psi$ and $v$ were arbitrary. ∎
@@
Since the map $B:V\rightarrow Q$ is surjective, it follows from the closed range Theorem, see e.g. [23, Lemma A.40], that there exists a constant $\beta>0$ with
@@
Combining 77 with 78 finishes the proof. ∎
@@
where $P_{W_{h}}$ is the $L^{2}$-projection operator onto $W_{h}$ 63.
@@
The following result is helpful to show that the best approximation inequalities in Theorem 3.4 and Lemmas 3.5 and 3.6 are stable as $h\to 0$. Under stronger (standard) assumptions, we can also derive error convergence rates in $\ell$ and $h$.
Lemma 3.7.
Let $\psi=\psi^{*}:=\psi_{u^{*}}$, $\|\psi_{h}^{0}\|_{L^{\infty}(\Omega_{d};\mathbb{R}^{m})}\lesssim 1$ and let Assumptions 1 and 2 hold. Assume that $\|u^{*}-\mathcal{E}_{h}u^{*}\|_{V}\rightarrow 0$ as $h\rightarrow 0$. Then
(80) $(\mathcal{D}^{*}(\psi_{h}^{0},\psi^{*}),1)_{\Omega_{d}}\leq C_{\mathrm{stab}},$
where $C_{\rm{stab}}$ is independent of the mesh size $h$ and iteration count $\ell$.
Proof.
Observe that $\nabla\mathcal{R}^{*}(\psi^{*})=B\,\mathcal{E}_{h}u^{*}$ in $\Omega_{d}$ by 64 and 26. In addition, using 29, we deduce that
$\mathcal{R}^{*}(\psi^{*})=\psi^{*}(B\mathcal{E}_{h}u^{*})-(\mathcal{R}\circ B)(\mathcal{E}_{h}u^{*}).$
Using these observations and the definition of $\mathcal{D}^{*}$ 34, we can rewrite
(81) $\int_{\Omega_{d}}\mathcal{D}^{*}(\psi_{h}^{0},\psi^{*})\operatorname{d}\!\mathcal{H}_{d}$ $=\int_{\Omega_{d}}(\mathcal{R}^{*}(\psi_{h}^{0})-\mathcal{R}^{*}(\psi^{*})-\nabla\mathcal{R}^{*}(\psi^{*})(\psi_{h}^{0}-\psi^{*}))\operatorname{d}\!\mathcal{H}_{d}$
$=\int_{\Omega_{d}}\mathcal{R}^{*}(\psi_{h}^{0})\operatorname{d}\!\mathcal{H}_{d}+\int_{\Omega_{d}}(\mathcal{R}\circ B)(\mathcal{E}_{h}u^{*})\operatorname{d}\!\mathcal{H}_{d}-\int_{\Omega_{d}}(B\mathcal{E}_{h}u^{*})\psi_{h}^{0}\operatorname{d}\!\mathcal{H}_{d}.$
The first term is bounded since $\psi_{h}^{0}$ is uniformly bounded and $\mathcal{R}^{\ast}$ is continuous on $L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. For the second term, we note that $\mathcal{E}_{h}u^{*}\in K$ and by assumption $\|u^{*}-\mathcal{E}_{h}u^{*}\|_{V}\rightarrow 0$ as $h\rightarrow 0$. From the continuity of $B$ and the assumed continuity of $\mathcal{R}$ stated in Assumption 1, we have that
(82) $\int_{\Omega_{d}}(\mathcal{R}\circ B)(\mathcal{E}_{h}u^{*})\operatorname{d}\!\mathcal{H}_{d}\rightarrow\int_{\Omega_{d}}(\mathcal{R}\circ B)(u^{*})\operatorname{d}\!\mathcal{H}_{d}.$
This provides a uniform bound on the second term in 81. For the last term in 81, a uniform bound is obtained by using the Cauchy–Schwarz inequality, the observation that $\|Bu^{*}-B\mathcal{E}_{h}u^{*}\|_{L^{2}(\Omega_{d})}\rightarrow 0$ as $h\rightarrow 0$, which follows from the continuity of the map $B$, and the assumption that $\|\psi_{h}^{0}\|_{L^{\infty}(\Omega_{d};\mathbb{R}^{m})}\lesssim 1$. This concludes the proof. ∎
@@ 3.5. Convergence rates and mesh-independence
We now derive abstract error convergence rates from Theorem 3.4 and Lemmas 3.5 and 3.6. For simplicity, we focus on the following spaces $V\subset H^{1}(\Omega)$ and $V\subset H^{1}(\Omega;\mathbb{R}^{n})$. Hereafter, we do not explicitly differentiate between scalar and vector-valued spaces, using the notation for the scalar space while noting that the same results hold, mutatis mutandis, for vector-valued spaces. To derive error rates, we require the following quasi-interpolation assumption on the stability and approximation of the Fortin and enriching operators. The decomposition of the enriching map given in 84 reflects the constructions used in Sections 4 and 5, below.
@@
Since $E^{\prime}(u^{*})\in H^{r-1}(\Omega)$, we bound the last term in 69 by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}(\|u_{h}^{\ell}-v\|_{H^{1-r}(\Omega)}+\|u^{*}-\mathcal{U}_{h}(\psi)\|_{H^{1-r}(\Omega)})$. Thus, we obtain
(87) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{H^{1}(\Omega)}^{2}\leq\inf_{\psi\in\Psi_{h},v\in K}\bigg{(}\frac{(\mathcal{D}^{*}(\psi_{h}^{0},\psi),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|u^{\ast}-\mathcal{U}_{h}(\psi)\|_{H^{1}(\Omega)}^{2}\\ \|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\big{(}\|u_{h}^{\ell}-v\|_{H^{1-r}(\Omega)}+\|u^{*}-\mathcal{U}_{h}(\psi)\|_{H^{1-r}(\Omega)}\big{)}\bigg{)}.$
We now proceed by bounding each term in 87. We select $\psi=\psi^{*}:=\psi_{u^{*}}$. Note that from 85b, 85c, and a triangle inequality, we can conclude that $\|u^{*}-\mathcal{E}_{h}u^{*}\|_{V}\rightarrow 0$ as $h\rightarrow 0$. Thus, Lemma 3.7 applies and $(\mathcal{D}^{*}(\psi_{h}^{0},\psi^{*}),1)_{\Omega_{d}}\lesssim C_{\mathrm{stab}}$.
Proceeding, we choose to bound $\|u^{*}-\mathcal{U}_{h}(\psi^{*})\|_{H^{1}(\Omega)}$. Using the definition of $\mathcal{U}_{h}$ 65, the triangle inequality, and 85, we obtain that
(88) $\|u^{*}-\mathcal{U}_{h}(\psi^{*})\|_{L^{2}(\Omega)}$ $\leq\|u^{*}-\Pi_{h}u^{*}\|_{L^{2}(\Omega)}+\|\Pi_{h}(u^{*}-\mathcal{E}_{h}u^{*})\|_{L^{2}(\Omega)}$
$\lesssim\|u^{*}-\Pi_{h}u^{*}\|_{L^{2}(\Omega)}+\|u^{*}-\mathcal{E}_{h}u^{*}\|_{L^{2}(\Omega)}+h\|\nabla(u^{*}-\mathcal{E}_{h}u^{*})\|_{L^{2}(\Omega)}$
$\lesssim h^{1+s}(\|u^{*}\|_{H^{1+s}(\Omega)}+1).$
A similar argument shows that the same bound holds on $h\|\nabla(u^{*}-\mathcal{U}_{h}(\psi^{*}))\|_{L^{2}(\Omega)}$. Therefore,
(89) $\|u^{*}-\mathcal{U}_{h}(\psi^{*})\|^{2}_{H^{1}(\Omega)}\lesssim h^{2s}(\|u^{*}\|_{H^{1+s}(\Omega)}+1)^{2}.$
To handle the $\|u^{*}-\mathcal{U}_{h}(\psi^{*})\|_{H^{1-r}(\Omega)}$ term in 87, we first note that from space interpolation between $L^{2}(\Omega)$ and $H^{1}(\Omega)$ (see, e.g., [9, Chapter 14, Proposition 14.1.5]) and 85a, we obtain that
@@
In addition, owing to the Gagliardo–Nirenberg inequality [12, Theorem 1] and Young’s inequality ($a^{\alpha}b^{\beta}\leq\alpha a+\beta b$ for $\alpha,\beta\geq 0$ and $\alpha+\beta=1$), we have that
(91) $\|\Pi_{h}w\|_{H^{t}(\Omega)}\lesssim\|\Pi_{h}w\|_{L^{2}(\Omega)}^{1-t}\|\Pi_{h}w\|_{H^{1}(\Omega)}^{t}\\ =(h^{-t}\|\Pi_{h}w\|_{L^{2}(\Omega)})^{1-t}(h^{1-t}\|\Pi_{h}w\|_{H^{1}(\Omega)})^{t}\lesssim h^{-t}\|\Pi_{h}w\|_{L^{2}(\Omega)}+h^{1-t}\|\Pi_{h}w\|_{H^{1}(\Omega)}.$
Therefore, applying the triangle inequality, the above estimates with 85c and $t=1-r$, 83, and 85, we obtain that
(92) $\|u^{*}-\mathcal{U}_{h}(\psi^{*})\|_{H^{1-r}(\Omega)}$ $\leq\|u^{*}-\Pi_{h}u^{*}\|_{H^{1-r}(\Omega)}+\|\Pi_{h}(u^{*}-\mathcal{E}_{h}u^{*})\|_{H^{1-r}(\Omega)}$
$\lesssim h^{r+s}|u^{*}|_{H^{1+s}(\Omega)}+h^{-1+r}\|\Pi_{h}(u^{*}-\mathcal{E}_{h}u^{*})\|_{L^{2}(\Omega)}+h^{r}\|\Pi_{h}(u^{*}-\mathcal{E}_{h}u^{*})\|_{H^{1}(\Omega)}$
$\lesssim h^{r+s}|u^{*}|_{H^{1+s}(\Omega)}+h^{-1+r}\|u^{*}-\mathcal{E}_{h}u^{*}\|_{L^{2}(\Omega)}+h^{r}\|u^{*}-\mathcal{E}_{h}u^{*}\|_{H^{1}(\Omega)}$
$\lesssim h^{r+s}|u^{*}|_{H^{1+s}(\Omega)}.$
It remains to estimate the term in 87 involving $\|u_{h}^{\ell}-v\|_{H^{1-r}(\Omega)}$. To this end, we choose $v=\mathcal{E}_{h}(u_{h}^{\ell})$. This is a valid choice since $u_{h}^{\ell}\in K_{h}$ by 10b and $\mathcal{E}_{h}u_{h}^{\ell}\in K$ by Assumption 2. With the help of 85, we now formulate the following estimate:
@@
where $c$ and $\tilde{c}$ are mesh-independent constants. Incorporating 80, 89, 92, and 93 into 87 yields the required bound on $\|u_{h}^{\ell}-u^{*}\|_{H^{1}(\Omega)}$.
The estimate on $\|\lambda^{*}-\lambda_{h}^{\ell}\|_{H^{-1}(\Omega)}$ follows from Lemma 3.5. In particular, the first term in 76 is bounded by:
@@
using 90 for the second inequality. ∎
The final result of this section follows from the fact that the approximation error in 86 is controlled by independent optimization and discretization error terms, each depending only on $\ell$ and $h$, respectively.
@@ 3.6. Discussion of results
Theorems 3.4 and 3.8 rigorously demonstrate three important features of the PG method, Algorithm 1, for the first time. To date, these features have only been observed numerically [36, 28, 21, 40].
  • First, the PG iterates $u^{\ell}_{h}$ converge to the true solution $u^{\ast}$ as $\ell\to\infty$ and $h\to 0$, even with bounded proximity parameters, $\alpha_{k}\leq\textrm{const.}$. This stands in contrast to penalty [30, Chapter 1.7] and interior point methods [38, Chapter 19] — and even augmented Lagrangian methods in infinite-dimensional spaces [35, 3] — which all require taking a relaxation parameter to a singular limit.
@@
  • Finally, PG can be applied without modification to problems with low-regularity solutions or multipliers. Indeed, Theorem 3.8 provides sufficient conditions on $\Pi_{h}$ and $\mathcal{E}_{h}$ to obtain optimal convergence rates in $h$ depending naturally on the solution $u^{\ast}$ and Fréchet Derivative $E^{\prime}(u^{\ast})$. The following two sections show that these abstract conditions are checkable in practice. It is well-known that penalty methods can also be applied without modification in the low-regularity setting [33]. However, to the best of our knowledge, the existing theory for discretized penalty methods requires $u^{*}\in H^{1+s}(\Omega)$ and $s\in(1/2,1]$ when applied to the obstacle and Signorini problems. See [44, 31] for the obstacle problem and [16, Section 6] for the Signorini problem. Notably, the PG results hold for any $s\in(0,1]$.

### v2

@@ 3.4. Best approximation error estimates
(65) $E(u_{h}^{\ell})-E(u)+\frac{1}{2}\nu\|u-\overline{u}_{h}^{\ell}\|_{V}^{2}+\frac{(\mathcal{D}(Bu,o_{h}^{\ell}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}\\ \leq\frac{(\mathcal{D}(Bu,o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\langle E^{\prime}(\overline{u}_{h}^{\ell})-E^{\prime}(u),\Pi_{h}u-u\rangle+\langle E^{\prime}(u),\Pi_{h}u-u\rangle.$
Proceeding, we use the continuity of $a$ (8) and Young’s inequality to bound
$\langle E^{\prime}(\overline{u}_{h}^{\ell})-E^{\prime}(u),\Pi_{h}u-u\rangle\leq M\|\overline{u}_{h}^{\ell}-u\|_{V}\|\Pi_{h}u-u\|_{V}\leq\frac{\nu}{2}\|\overline{u}_{h}^{\ell}-u\|_{V}^{2}+\frac{M^{2}}{2\nu}\|\Pi_{h}u-u\|_{V}^{2}.$
Combining the above estimates yields
(66) $E(u_{h}^{\ell})-E(u)+\frac{(\mathcal{D}(Bu,o_{h}^{\ell}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}\\ \leq\frac{(\mathcal{D}(Bu,o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u-u\|^{2}_{V}+\langle E^{\prime}(u),\Pi_{h}u-u\rangle.$
At this stage, we select $u=u^{*}$ and note that for any $v\in K$, 59 and 9 provide the bound
@@
Thus, adding $\langle E^{\prime}(u^{\ast}),v-u_{h}^{\ell}\rangle$ to both sides of 66 and recalling that $Bu^{*}=o^{*}$, we arrive at
(67) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}+\frac{(\mathcal{D}(o^{*},o_{h}^{\ell}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}\leq\frac{(\mathcal{D}(o^{*},o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u^{*}-u^{*}\|^{2}_{V}\\ +|\langle E^{\prime}(u^{*}),\Pi_{h}u^{*}-u^{*}\rangle|+|\langle E^{\prime}(u^{*}),v-u_{h}^{\ell}\rangle|.$
The result follows because the choice $v$ was arbitrary and $\mathcal{D}(o^{*},o_{h}^{\ell})\geq 0$. ∎
@@
Since the map $B:V\rightarrow Q$ is surjective, it follows from the closed range Theorem, see e.g. [22, Lemma A.40], that there exists a constant $\beta>0$ with
@@
Combining 69 with 70 finishes the proof. ∎
@@
where $P_{W_{h}}$ is the $L^{2}$-projection operator onto $W_{h}$ 75.
@@ 3.5. Convergence rates and mesh-independence
We now derive abstract error convergence rates from Theorem 3.3 and Lemmas 3.4 and 3.5. For simplicity, we focus on the following spaces $V\subset H^{1}(\Omega)$ and $V\subset H^{1}(\Omega;\mathbb{R}^{n})$. Hereafter, we do not explicitly differentiate between scalar and vector-valued spaces, using the notation for the scalar space while noting that the same results hold, mutatis mutandis, for vector-valued spaces. To derive error rates, we require the construction of an enriching map that allows us to control the last term in (58). In turn, we make the following general assumption, which is critical to defining this map. First, we motivate this assumption with a remark.
Remark 1 (Motivating Assumption 2).
As expected from the analysis of nonconforming discretizations of obstacle problems [5, Section 5.2.1], one then requires an operator $\mathcal{E}_{h}:K_{h}\rightarrow K$ with suitable approximation properties. See Section 4.3 for an explicit construction of the enriching map for $K_{h}$ defined in 72.
Assumption 2.
(73) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{E}_{h}:K\cup K_{h}\rightarrow K$},
where
The set $K_{h}$ contains functions $u_{h}\in V_{h}$ such that $Bu_{h}\in W_{h}$ is the $L^{2}$-projection of a constrained observable $o\in\mathcal{O}$. It is constructed to ensure that $u_{h}^{k}\in K_{h}$ for all iterations $k$; see 10b. To derive error rates, we require the following quasi-interpolation assumption on the stability and approximation of the Fortin and enriching operators. The decomposition of the enriching map given in 77 reflects the constructions used in Sections 4 and 5, below.
@@
Since $E^{\prime}(u^{*})\in H^{r-1}(\Omega)$, we further bound the last two terms in 58 by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}$ and by $\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$ respectively. Thus, we obtain
(80) $\frac{\nu}{2}\|u^{*}-u_{h}^{\ell}\|_{V}^{2}\leq\frac{(\mathcal{D}(o^{*},o_{h}^{0}),1)_{\Omega_{d}}}{\sum_{k=1}^{\ell}\alpha_{k}}+\frac{M^{2}}{2\nu}\|\Pi_{h}u^{*}-u^{*}\|^{2}_{H^{1}(\Omega)}\\ +\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}+\inf_{v\in K}\|E^{\prime}(u^{*})\|_{H^{r-1}(\Omega)}\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}.$
We now proceed by bounding each term in 80. Recalling that $o_{h}^{0}=\nabla\mathcal{R}^{*}(\psi_{h}^{0})$, $\|\psi_{h}^{0}\|_{L^{\infty}(\Omega_{d};\mathbb{R}^{m})}\lesssim 1$, and the definition of $\mathcal{D}$ (33), it readily follows that
(81) $\mathcal{D}(o^{*},o_{h}^{0})\leq C_{\mathrm{stab}},$
for some constant $C_{\mathrm{stab}}$ independent of $h,\ell$ and $\alpha_{k}$ since $\mathcal{R}$ is continuous on $\mathcal{O}$ and $\nabla\mathcal{R}^{*}$ is continuous on $L^{\infty}(\Omega_{d};\mathbb{R}^{m})$.
Proceeding, Assumption 3 directly provides the bound
(82) $\|u^{*}-\Pi_{h}u^{*}\|_{H^{1}(\Omega)}^{2}\lesssim h^{2s}|u^{*}|^{2}_{H^{1+s}(\Omega)}.$
To handle the $\|u^{*}-\Pi_{h}u^{*}\|_{H^{1-r}(\Omega)}$ term in 80, we first note that from space interpolation between $L^{2}(\Omega)$ and $H^{1}(\Omega)$ (see, e.g., [8, Chapter 14, Proposition 14.1.5]) and 78a, we obtain that for any $w\in H^{1+s}(\Omega)$,
@@
Therefore, we obtain that
(84) $\|\Pi_{h}u^{*}-u^{*}\|_{H^{1-r}(\Omega)}\lesssim h^{r+s}|u^{*}|_{H^{1+s}(\Omega)}.$
It remains to estimate the term in 80 involving $\|v-u_{h}^{\ell}\|_{H^{1-r}(\Omega)}$. To this end, we choose $v=\mathcal{E}_{h}(u_{h}^{\ell})$. This is a valid choice since $u_{h}^{\ell}\in K_{h}$ by 10b and $\mathcal{E}_{h}u_{h}^{\ell}\in K$ by Assumption 2. With the help of 78 and (83), we now formulate the following estimate:
@@
where $c$ and $\tilde{c}$ are mesh-independent constants. Incorporating 81, 82, 84, and 85 into 80 yields the required bound on $\|u_{h}^{\ell}-u^{*}\|_{H^{1}(\Omega)}$.
The estimate on $\|\lambda^{*}-\lambda_{h}^{\ell}\|_{H^{-1}(\Omega)}$ follows from Lemma 3.4. In particular, the first term in 68 is bounded by:
@@
using 83 for the second inequality. ∎
The final result of this section follows from the fact that the approximation error in 79 is controlled by independent optimization and discretization error terms, each depending only on $\ell$ and $h$, respectively.
@@ 3.6. Discussion of results
Theorems 3.3 and 3.6 rigorously demonstrate three important features of the PG method, Algorithm 1, for the first time. To date, these features have only been observed numerically [36, 28, 19, 40].
  • First, the PG iterates $u^{\ell}_{h}$ converge to the true solution $u^{\ast}$ as $\ell\to\infty$ and $h\to 0$, even with bounded proximity parameters, $\alpha_{k}\leq\textrm{const.}$ This stands in contrast to penalty [30, Chapter 1.7] and interior point methods [38, Chapter 19] — and even augmented Lagrangian methods in infinite-dimensional spaces [35, 3] — which all require taking a relaxation parameter to a singular limit.
@@
  • Finally, PG can be applied without modification to problems with low-regularity solutions or multipliers. Indeed, Theorem 3.6 provides sufficient conditions on $\Pi_{h}$ and $\mathcal{E}_{h}$ to obtain optimal convergence rates in $h$ depending naturally on the solution $u^{\ast}$ and Fréchet Derivative $E^{\prime}(u^{\ast})$. The following two sections show that these abstract conditions are checkable in practice. It is well-known that penalty methods can also be applied without modification in the low-regularity setting [33]. However, to the best of our knowledge, the existing theory for discretized penalty methods requires $u^{*}\in H^{1+s}(\Omega)$ and $s\in(1/2,1]$ when applied to the obstacle and Signorini problems. See [44, 31] for the obstacle problem and [14, Section 6] for the Signorini problem. Notably, the PG results hold for any $s\in(0,1]$.